Prev Next

SW&IT Meeting Minutes 2019-07-08

Attendees: Andrew, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang


Labor Reporting

Report auto-generated every Monday at 6.30AM, Do not edit this table, use entries in smartsheets.

LR20190701 - Bec, Matthieu 51.5
LR20190701 - Colson, Andrew 16.0
LR20190701 - Edmeades, David 32.0
LR20190701 - Sola, Igor 0.0
LR20190701 - Kubanek, Petr 24.0
LR20190701 - Hooper, Stephen 64.0
LR20190701 - Zhang, Yang 24.0
LR20190701 - Zhang, Xianyu 25.0
LR20190701 - Holmberg, Carl 0.0

News from management

  • mirror cover report being reviewed.

Previous Action Items

  • POI/POR form for observing workstation upgrade (SH)
  • update obs ws upgrade working document (IS, XZ, YZ, MB)
  • review POR form for colocation (MB)
  • describe TCS 2019A release strategy: split ICE as separate upgrade? test plan for different IIF clients, reachout to LN, PEPSI, etc.) (PK)
  • describe AO telemetry and reduction strategy, need for storage (XZ)
  • UITS coloc, use the POI

  • ports 4000 and 5000 on (sciarc) test machines
  • 'airflow' email group
  • polycom for italian partners
  • Updates to MS terms of use ?
  • supportscripts automated git backup, Permissions 0644 for '/etc/ssh/swgroup-github.key' are too open.

Open topics

  • TCS (and dependencies) does not compile with newer compilers (GCC 8.x). I encountered the problem while trying to compile on recent Ubuntu. I have some code working, shall we create a global effort to make sure everything compiles with the current GCC?

Team reports


  • OC, WPC, Management
  • complete PEs
  • GMT M1 PDR written feedback report
  • misc prep for CH
  • contibutions to obs-ws-up POI
  • review UITS coloc POR
  • review PG's remote sensors and API
  • CH induction
  • misc. planning activities
  • 3 weeks vacation


  • Fixed AUP link in registration form
  • Matlab licensing update
    • Test stand: done
    • Jet: problem with old software version and being unable to generate licenses that old. Opened ticket with Mathworks. Kicked back due to me not being a license admin, talking to the UA administrators.
  • Wireless sensor setup and testing
  • Zoom migration complete
    • $350 refund and no charges going forward
    • Tested swgroup meeting with success (moot at the point of reading this)
    • Note that all new meeting links now start with "". "" links appear to still work.
  • VMHost2 networking
    • Down to one vSwitch
    • Console op to move management network pending
  • Safety training
  • POI form start
  • now working; was ownership of main program directory. Still no insight on what caused the initial failure.
  • Second Elliot workstation
  • POI work
  • Files.lbto last attempt to know what happened
  • VMHost 2 management network swap
  • Created airflow group per CV request
  • New Zoom role accounts created: lbto-zoom[3,4,5]


  • Still reworking the Python LBC skyflats wiki program.
  • Getting back to AO-Cravat web code.
  • MODS IT#7736, Rick and Olga remapped the hardware limit. Software limit to be updated sometime later in the code. Also needs to be updated in OT. An interface control document is planned.
  • MODS IT#7764, verified the IIF interface info. coordinating with Olga for further needs.
  • Need to organize the sciops-lbcfpia/lbcwrs github repository.
  • Reviewing the sciops-lbttools repository as well.
  • Discussed with Olga and Barry regarding the remote room machine and obs machine upgrade. Gathered some local programs need to be configured and installed on remote room machines. Still onging, more to be discussed next.


  • safety training
  • pms-lasair - web interface down. Can ping but can't telnet.
  • ARGOS OBSWUP meeting in August
  • test FACSUM on remote obs NUC
    • IT 6405 - weather charts updating irregularly
    • IT 7705 - values updating irregularly
  • bring pms-lasair back online and telemetry


  • IT 7746 - IRC verification that it works without ACM
  • TCS under new GCC )8.2)


  • kickstart workstations
  • created RPMs for filesystem mount points
  • VPN accounts for iLocater people
  • iLocater DNS changes
  • add accounts for new employees
  • POI (briefly)
  • Wiese computer issues
  • RPMs for missing lbto.path (perl) and xephem (F30 issue)
  • rm507 test (why doesn't systemd work with the NFS mounts)
  • archive things.


Last week
  • OT
    • Still working on position editor submenus. Michelle and I went thoroughly through all the items, about 80% done. There stil a number of things we'd like to get done, such as removing some submenus, or moving the submenu behaviour to the parent menu.
    • LBC subwindow chips, Chip4 is an outlier and requires special coding, which I hate and might be flaky trying to find a way to keep existing code, but with minimal changes so it doesn't disrupt existing overall logic.
  • LUCI
    • Working with IT to come up with a satisfactory LUCI script and directory structure to be used in the observer machines.

This week
  • OT
    • #IT7441 subwindowing for LBC chips
    • #IT7074, #IT7192 Issues related to ARGOS and ARGOS guide stars and also the way OT handles errors.


Last week
  • 4D port test
  • process the data to determin jumping actutor in May and Jun 2019
  • Safety training for few hours
  • TCS Ice interface info with MPIA (LN and ARGOS)
This week
  • ASM summer shutdown plan
  • Discuss jumping actutor results with Guido
  • Save ASM currents in TN?
  • Build SOUL AdSec simulator
  • Safety training?
Topic revision: r21 - 15 Jul 2019, IgorSola
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback